OpenStudy (anonymous):

Noel currently has a glance of 2,300.37 in an account he has held for 34 years. He opened the account with an initial deposit of $874. What is the simple interest rate on the account? I keep getting an off number /:

OpenStudy (deepolisnoob):

The equation is Future value= Present value (1+interest rate/times compounded annually)^years*times compounded annually

OpenStudy (anonymous):

\[A = P(1 + rt)\]=\[2,300.37=874(1+34r)\] solve for \[r\]

OpenStudy (anonymous):

\[r=0.048 \]=4.8%

OpenStudy (anonymous):

Thank you! I was dividing my years by 12, lol.

OpenStudy (anonymous):

I mean multiplying them

OpenStudy (anonymous):

;)